OpenStudy (anonymous):

Algebra 1 help! Fan and medal!!

OpenStudy (anonymous):

What do you need?

OpenStudy (anonymous):

1. Esmeralda and Heinz are working to graph a polynomial function, f(x). Esmeralda says that the third-degree polynomial has four intercepts. Heinz argues that the function only crosses the x-axis three times. Is there a way for them both to be correct? Explain your answer.

OpenStudy (anonymous):

That is the first question. There's 4

OpenStudy (anonymous):

@ANA789 do you think you can help?

OpenStudy (anonymous):

Please guys I need help!

OpenStudy (anonymous):

@GeniousCreation please I beg of you, can you help me?

OpenStudy (anonymous):

@mely1014 Im sorry Im not very good at equations like this

OpenStudy (anonymous):

@ANA789 well do you know someone who can help me?

OpenStudy (anonymous):

@Robert136

OpenStudy (anonymous):

@xapproachesinfinity can you help??

OpenStudy (xapproachesinfinity):

actually no! according to the fundamental theorem of algebra a third degree if its solutions are all real then i will cross the x axis three times if not then it will have at least one real solution 4 intercepts means 4 real solutions which is not true for a third degree polynomial

OpenStudy (anonymous):

@xapproachesinfinity OK so I would put it like this- No because a third degree has to have all real solutions for it to pass the x-axis three times.

OpenStudy (anonymous):

@Flvs_student21 I used to go to flvs, think you can help me with the other questions?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

2. Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). • g(x) = x3 – x2 – 4x + 4 • g(x) = x3 + 2x2 – 9x – 18 • g(x) = x3 – 3x2 – 4x + 12 • g(x) = x3 + 2x2 – 25x – 50 • g(x) = 2x3 + 14x2 – 2x – 14 This is the second question

OpenStudy (wil476003):

I found this this for you I am not this far in math yet but it is somebody helping someone else out with the problems and it made since to me I hope this helps I'm sorry I couldn't really help u http://openstudy.com/study#/updates/5349df3ae4b0dfc5ac7179f9

OpenStudy (anonymous):

@wil476003 it helps in the first 3. Thank you! I still need the last one though but thanks

OpenStudy (wil476003):

Yurp I'm sorry if I new it I would help you solve it all

OpenStudy (anonymous):

@wil476003 this is more than enough. Thanks again:)

OpenStudy (xapproachesinfinity):

No that is not what i said! i said that a third degree can have a maximum of 3 real solution it can't have more than that

OpenStudy (anonymous):

@xapproachesinfinity calm down haha. It's what I understood